A circle is divided into 12 equal pieces shade 1/3 of it
Answer by addingup(3677)

Well, if you don't have the circle with the slices you need to draw one. Then,
12*1/3 = (12*1)/3 = 12/3 divide and then you'll know how many slices you have to shade.